Ex 6.1 Class 8 Maths Question 6.
Using the given pattern, find the missing numbers.
12 + 22 + 22 = 32
22 + 32 + 62 = 72
32 + 42 + 122 = 132
42 + 52 + ….2 = 212
52 + ….2 + 302 = 312
62 + 72 + …..2 = ……2

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Solution: According to the given pattern, we have 42 + 52 + 202 = 212 52 + 62 + 302 = 312 62 + 72 + 422 = 432